# A POD-TANN approach for the multiscale modeling of materials and macroelement derivation in geomechanics

### v2

@@ Abstract
The methodology leverages POD to extract macroscopic Internal State Variables (ISVs) from microscopic state information, thereby enriching the macroscopic state description used to train an energy potential network within the TANN framework. The thermodynamic consistency provided by TANN, combined with the hierarchical nature of POD, allows for accurate modeling of complex, non-linear material behavior and reliable macroscopic geomechanical systems responses.
The effectiveness of this approach is validated through applications of increasing complexity, demonstrating its capability to handle various material behaviors and microstructural topologies. These applications include the homogenization of continuous inelastic representative unit cells (RUCs) and the derivation of a macroelement for a geotechnical system involving a monopile in a clay layer subjected to horizontal loading.
The results indicate that the proposed POD-TANN methodology not only achieves high accuracy in reproducing stress-strain responses, but also significantly reduces computational costs, making it a practical tool for the multiscale modeling of heterogeneous inelastic systems, and the efficient derivation of macroelements for complex geomechanical problems.
@@ 1 Introduction
There are two main challenges in establishing a machine learning (ML)-based homogenization procedure. First, the obtained model at the macroscale must comply with basic physical laws, a general requirement not specific to ML techniques adopted to derive it. This necessitates incorporating such laws into the structure of the ML tools. Recently, neural network-based approaches have been proposed that incorporate basic physical laws into the model structure by design (see [1], [2], [3], [4], among others). Second, the multi-scale problem must account for the properties and responses of the constituents at the microscopic level. Examples of ML-based strategies for accelerating multi-scale analyses can be found in [5], [6].
@@
In this work, POD is applied to a set of microscopic state variables to extract the internal state variables at the macroscopic scale as a result of dimensionality reduction. Unlike other techniques, POD offers an unsupervised identification methodology that is hierarchical, preserving maximal information in the reduced space given the dimension of the embedding. This feature is exploited in the paper to select a sufficiently large space of ISVs for training the energy network, a neural network used to learn the macroscopic (Helmholtz) free energy of the composite.
The hierarchical nature of POD, along with its robustness and efficiency, significantly reduces the computational cost by decoupling the ISV identification from the training of the energy network.
@@ 4 Applications
This section will examine three applications of the proposed approach. Sections 4.1 and 4.2 will focus on learning the homogenized behavior of continuum microstructured periodic unit cells exhibiting inelastic behavior. Then, Section 4.3 will detail the data-driven derivation of a macroelement for the horizontal response of a monopile in a clayey layer.
@@ 4.3.3 Inference of the TANN
4.3.4 Discussion on the POD-TANN macroelement calibration
@@
An important consideration in this context is the number of POD modes used in the model, as this directly influences the dimensionality of the internal state variable vector. The selection of POD modes is crucial, as it determines how well the reduced-order model can capture the essential features of the system. For TL to be effective across different scenarios, the selected number of POD modes should be representative across a range of macro-model parameters, including variations in geometry, material properties, and boundary conditions, as stated before. This requirement would ensures that the fixed number of POD modes (i.e. of ISV vector components) remains unchanged and sufficient for new scenarios, thereby enabling the successful application of TL, without TANN’s hyperparameters changes (e.g., the number of inputs, and consequently the number and density of hidden layers).
@@ 5 Discussion and conclusions
A new POD-TANN methodology has been presented, which incorporates Proper Orthogonal Decomposition into the training workflow of Thermodynamics-based Artificial Neural Networks [7], for modeling the macroscopic behavior of inelastic systems with microstructure.
The POD is applied to datasets of snapshots of internal coordinates, i.e., to the set of state variables apt to describe the processes taking place at the microscale of the investigated heterogeneous, inelastic materials. The method places at its core the hierarchical structure of the modes obtained via POD, which guarantees the maximum representativeness of the information contained in the obtained first $N$-th modes. The consequent hierarchical sorting of the ISV vector’s components allows decoupling the dimensionality reduction procedure from the training of the TANN network. This procedure offers great versatility. Indeed, the number of the ISV’s components can be selected based on the degree of accuracy depending on the applications.
One key benefit of this approach is the possibility of including microscopic information at the macroscale through a linear dimensionality reduction technique that is straightforward to implement, fast, and robust. The linear mapping is guaranteed to be invertible, making the reconstruction of the microscopic fields always possible. This feature is very compelling since it provides the direct and inverse mapping to perform homogenization at the macroscale and the reconstruction of the fields at the microscale, respectively. However, this feature is influenced by the number of POD modes considered. When only the macroscopic behavior is of interest, a reduced number of modes can be selected, aiming only at training the TANN network within a desired accuracy. If the microscopic fields are the primary objective, the whole procedure can be used as a reduced order modeling tool, and the number of POD modes can be selected to reduce the reconstruction error of the microscopic fields below a given tolerance.
The effectiveness of the methodology is demonstrated through three applications of increasing complexity. These include the homogenization of two inelastic RUCs with continuous microstructure, featuring inclusions of increasing complexity. Furthermore, a complex three-dimensional geotechnical model of a monopile in a clay layer subjected to horizontal loading is studied to demonstrate the applicability of the procedure for the derivation of macroelements.
In all cases, the trained network was able to reproduce accurately the stress-strain (or force-displacement) paths both within and outside the training range. Furthermore, the reconstructed microscopic fields are presented, specifying the reconstruction error as a function of the selected number of POD modes.
Throughout the paper, a loss function that excludes energy terms was used for training the energy network (see Section 3.3). The effectiveness of this function is demonstrated in the application section. By omitting energy terms, the procedure becomes applicable to data from various sources, such as experimental data, where energy measurements are challenging to obtain.
Despite the linear nature of the dimensionality reduction method, the resulting compression ratios - defined as defined as $CR=1-\text{dim}(\boldsymbol{\mathcal{Z}})/\text{dim}(\boldsymbol{\xi})\%$ are highly satisfactory, consistently exceeding 90%. Notably, for RUCs with ellipsoidal and leaf-shaped inclusions, IC sets encompassing approximately 100000 degrees of freedom are reduced to an ISV vector with only 25 components, yielding a $CR$ of 99.97%. This high compression ratio significantly reduces the complexity of the hidden layers in the TANN energy network. In the examples considered, a network with a single hidden layer of 100 neurons was sufficient. In most cases, training converged rapidly, typically after 1000 epochs of optimization. The entire procedure requires approximately five minutes of runtime on standard 64-bit machines with 4 cores. Thus, the method is not only accurate but also highly efficient. The macroscopic stress predictions generated by this methodology have a mean absolute error on the order of 1e-4.
The proposed POD-TANN approach may also serve as a valuable preliminary analysis tool for materials with more complex microstructures. This is particularly relevant for materials exhibiting strong nonlinearities, such as those undergoing softening, where localization phenomena within the RUC can take place. In such cases, classical homogenization techniques, like Asymptotic Expansion Homogenization, typically fail to account for bifurcation or localization phenomena, limiting their applicability in the presence of these complex behaviors.
The challenge with strong nonlinearities lies in the absence of a general homogenization framework that can accurately handle such cases. While AEH provides a rigorous approach for homogenizing periodic microstructures, its formulation does not allow the inclusion of bifurcation phenomena. This limitation prevents its direct application in scenarios where localization occurs, as it is inherently restricted to smooth solutions within a standard continuum framework.
However, the TANN formulation, as presented here, is not tied exclusively to AEH or any specific homogenization technique. This opens the door for alternative frameworks to be considered. Importantly, the TANN methodology for homogeneous materials is not restricted to Cauchy continua, as demonstrated in previous work, [7]. Therefore, if a robust homogenization framework capable of accounting for bifurcation phenomena were developed, particularly one based on higher-order continuum theories, the proposed POD-TANN workflow could be applied.
Such a framework would enable the TANN to model more complex microstructures and material behaviors, including those influenced by localization and bifurcation, thereby broadening the scope of the approach.
5.1 Frontiers on the reconstruction task
An important aspect in the reconstruction of microscopic fields is the fulfillment of boundary conditions.
In the presented examples on periodic RUCs, the POD-TANN approach, while effective in reducing the dimensionality of the system and training efficient models, does not inherently guarantee the enforcement of periodic BCs in decoding. This limitation can introduce reconstruction errors, which depend on the quantity of interest and the various sources of error within the process.
More specifically, in cases where the primary goal is to reconstruct the strain field, the error in fulfilling periodic BCs tends to be directly related to the number of retained POD modes. Since the macroscopic strain is an input to the TANN (for strain-driven formulations), any mismatch between the predicted macroscopic strain and the microscopic strain field reconstruction will affect the overall periodicity. As such, the error will generally decrease as more POD modes are incorporated, ensuring a more accurate representation of the microscopic behavior. This relationship suggests the possibility of developing a POD mode selection criterion explicitly based on the error associated with periodic BC reconstruction. This could complement or even extend the energy-based selection criterion currently used for internal state variable (ISV) selection, allowing for a more tailored approach depending on whether accuracy in boundary condition fulfillment is prioritized. This goes to direction of applying PINNs losses for the resolution of BVPs, see, e.g., [43] or [44].
A key consideration here is that the reconstruction of microscopic stress fields is not directly governed by the macroscopic stress predictions from the TANN, as no explicit localization operator maps the macroscopic stress to its microscopic counterpart. Instead, the microscopic stress field is indirectly reconstructed through the strain field and the internal coordinates. Once the microscopic strain field and ISVs are accurately reconstructed, the stress can be computed using the microscopic constitutive laws of the material (always known in numerical examples, but possibly unknown for more general scenarios). This sequential process introduces additional error sources, particularly when periodic boundary conditions are involved, as any inaccuracy in the strain field reconstruction could propagate into the stress field. Therefore, ensuring the periodicity of the reconstructed strain field becomes crucial for minimizing errors in the subsequent stress field calculations.
Moreover, the complexity of accurately reconstructing the periodicity in fields such as strain or stress depends not only on the number of POD modes but also on the nature of the microstructure and the loading conditions. In systems with strong heterogeneities or non-linear behaviors, this error may grow, necessitating more sophisticated selection strategies or even hybrid approaches that combine linear reduction methods (such as POD) with non-linear techniques to capture more complex interactions at the micro-scale. By linking the number of POD modes directly to the fulfillment of boundary conditions, this could allow for a more nuanced trade-off between computational efficiency and reconstruction accuracy, especially in scenarios where enforcing periodicity is critical to the validity of the simulation.
In this view, a further advantage of the proposed procedure is its extensibility. The dimensionality reduction achieved through POD is linear and, in cases where the microstructure exhibits strong nonlinearities, may not yield a sufficiently small-dimensional ISV vector at the macroscale (irreducibility). To address this, nonlinear dimensionality reduction methods can be employed to further encode the POD-reduced set. For instance, POD can be used in conjunction with nonlinear autoencoders and SINDy, as demonstrated in [45] and [46]. Additionally, multiple local reduced-order bases (ROBs) can be utilized to manage non-reducible problems by decomposing them into multiple reducible ones, aided by physics-informed cluster analysis (see [47] and [48]).
5.1.1 Frontiers on the macroelement derivation and use
The developed technique demonstrates significant potential in reconstructing the structural field, particularly in the context of macroelement derivation. The ability to accurately reconstruct microscopic fields from macroscopic ISVs, as shown in the monopile example, underscores the practical utility of this methodology for engineering applications. By integrating POD and TANN, a robust and efficient framework is provided for both homogenization and localization, enhancing its applicability to complex geotechnical systems.
However, as highlighted in Section 4.3.4, the ”re-usability” of macroelements is currently limited due to the specificity of the computational setup, which is tailored to individual scenarios. This restriction necessitates retraining and reconfiguration when adapting to new conditions, such as changes in geometry, material properties, or boundary conditions. In this regard, Transfer Learning (TL) emerges as a promising avenue to overcome these limitations. By enabling the fine-tuning of pre-trained networks with smaller datasets, TL could extend the adaptability of macroelements across different scenarios, reducing the computational effort and enhancing the method’s scalability.
While the current framework is already effective for specific cases, the integration of TL presents a frontier for further expanding the flexibility and efficiency of the macroelement derivation process. This could enable the use of a generalized parametric macroelement model capable of adjusting to varying conditions, thereby broadening the application scope of the POD-TANN approach in geotechnical engineering.
Overall, the proposed methodology provides a powerful tool for researchers and practitioners in material science and engineering, offering thermodynamic consistency of the predicted system response based on TANN and the flexibility and power of model reduction offered by POD. Thus, the proposed POD-TANN approach shows high potential for modeling heterogeneous inelastic microstructured systems from data and predicting their response. For the class of applications shown, the results are very satisfactory and open the way for modeling even more complex scenarios.
@@ 6 Acknowledgment
The authors gratefully thank the funding of the European Research Council (ERC) under the Horizon 2020 research and innovation program of the European Union (Grant agreement ID 757848 CoQuake). In addition, the authors would like to thank Dr. Filippo Masi for their fruitful discussions, as well as the whole CoQuake research group.

### v3

@@ Abstract
The methodology leverages POD to extract macroscopic Internal State Variables from microscopic state information, thereby enriching the macroscopic state description used to train an energy potential network within the TANN framework. The thermodynamic consistency provided by TANN, combined with the hierarchical nature of POD, allows to reproduce complex, non-linear inelastic material behaviors as well as macroscopic geomechanical systems responses.
The approach is validated through applications of increasing complexity, demonstrating its capability to reproduce high-fidelity simulation data. The applications proposed include the homogenization of continuous inelastic representative unit cells and the derivation of a macroelement for a geotechnical system involving a monopile in a clay layer subjected to horizontal loading. Eventually, the projection operators directly obtained via POD, are exploit to easily reconstruct the microscopic fields.
The results indicate that the POD-TANN approach not only offers accuracy in reproducing the studied constitutive responses, but also reduces computational costs, making it a practical tool for the multiscale modeling of heterogeneous inelastic geomechanical systems.
@@ 1 Introduction
The complexity and variability of microscopic interactions, combined with the increasing availability of high-dimensional data, have motivated the exploration of machine learning (ML) techniques, which offer powerful tools for capturing intricate relationships that traditional methods may overlook.
There are two main challenges in establishing a ML-based homogenization procedure. Firstly, the obtained model at the macroscale must comply with basic physical laws, a general requirement not specific to ML techniques adopted to derive it. This necessitates incorporating such laws into the structure of the ML tools. Recently, neural network-based approaches have been proposed that incorporate basic physical laws into the model structure by design (see [1], [2], [3], [4], among others). Secondly, the multi-scale problem must account for the properties and responses of the constituents at the microscopic level. Examples of ML-based strategies for accelerating multi-scale analyses can be found in [5], [6].
@@
In this work, POD is applied to a set of microscopic state variables to extract the internal state variables at the macroscopic scale as a result of dimensionality reduction. Unlike other techniques, POD offers an unsupervised identification methodology that is hierarchical, preserving maximal information in the reduced space given the dimension of the embedding. This feature is exploited in the paper to select a sufficiently large space of ISVs for training the energy network, a neural network used to learn the macroscopic (Helmholtz) free energy of the composite. The hierarchical nature of POD, along with its robustness and efficiency, significantly reduces the computational cost by decoupling the ISV identification from the training of the energy network.
@@ 4 Applications
This section explores three applications of the proposed approach. Sections 4.1 and 4.2 focus on learning the homogenized behavior of periodic unit cells with inelastic microstructures, while Section 4.3 addresses the data-driven derivation of a macroelement for the horizontal response of a monopile embedded in a clay layer.
In these examples, both the macroscopic stress-strain (or force-displacement) response and the reconstruction of microscopic fields are examined, with particular emphasis on the influence of the number of retained POD modes for describing the components of the ISV vector. Accurately reconstructing microscopic fields, especially in systems using periodic Representative Unit Cells (RUCs), requires the fulfillment of boundary conditions. Although the POD-TANN approach effectively reduces system dimensionality and trains efficient models, it does not inherently enforce periodic boundary conditions during reconstruction, which may introduce errors. These errors depend on the quantity of interest and the various error sources involved, particularly in strain field reconstruction, where errors in fulfilling periodic BCs are typically related to the number of retained POD modes.
@@ 4.3.3 Inference of the TANN
5 Discussion and perspectives
5.1 Applicability to complex inelastic systems
The effectiveness of the proposed methodology is demonstrated through three applications of increasing complexity. These include the homogenization of two inelastic RUCs with continuous microstructures, featuring inclusions of increasing complexity. Additionally, a complex three-dimensional geotechnical model of a monopile embedded in a clay layer and subjected to horizontal loading is analyzed to showcase the procedure’s applicability in deriving macroelements.
At the core of the methodology is the use of POD to hierarchically select the components of the ISV vector from the set of ICs. Despite the linear nature of this dimensionality reduction, the resulting compression ratios in the proposed examples - function of the ratio between the dimensionality of the ISV vector and that of the IC set - are highly satisfactory, consistently exceeding 90%. In the RUC examples, IC sets comprising approximately 100000 degrees of freedom are reduced to an ISV vector with only 25 components, yielding a $CR$ of 99.97%. This high compression ratio significantly reduces the need for complex hidden layers of the TANN energy network. In the examples considered, a network with a single hidden layer of 100 neurons proved sufficient, and in most cases, training converged rapidly, typically after 1000 epochs of optimization. The entire procedure requires roughly five minutes of runtime on standard 64-bit machines with 4 cores. Thus, the method is not only accurate, with a mean absolute error of around 1e-4 in macroscopic stress predictions, but also highly efficient.
The proposed POD-TANN approach also may serve as a valuable tool for materials with more complex microstructures. This is particularly relevant for materials exhibiting strong heterogeneities and nonlinearities. To further test the potential of this approach in the presence of strong heterogeneities, an additional RUC is investigated in Appendix B, featuring continuous spatially correlated random fields of constitutive parameters. In this example we show again that number of the significant POD modes remains low and comparable with the examples presented in Sections 4.1 and 4.2. Morevoer, we report high accuracy under high compression rates.
A further advantage of the proposed procedure is its extensibility. The dimensionality reduction achieved through POD is linear and, in cases where the microstructure exhibits strong nonlinearities, may not yield a sufficiently small-dimensional ISV vector at the macroscale (irreducibility). To address this, nonlinear dimensionality reduction methods can be employed to further encode the POD-reduced set. For instance, POD can be used in conjunction with nonlinear autoencoders and SINDy, as demonstrated in [40] and [41]. Additionally, multiple local reduced-order bases (ROBs) can be utilized to manage non-reducible problems by decomposing them into multiple reducible ones, aided by physics-informed cluster analysis (see, e.g., [42] and [43]).
Classical homogenization techniques such as Asymptotic Expansion Homogenization (AEH) often fail to account for bifurcation or localization phenomena. While AEH provides a rigorous framework for homogenizing periodic microstructures, its formulation does not accommodate bifurcation phenomena, limiting its applicability in scenarios where localization occurs. AEH is inherently restricted to smooth solutions within a standard continuum framework. However, the TANN formulation presented here is not bound to AEH or any specific homogenization technique, making it adaptable to alternative frameworks. Notably, the TANN methodology for homogeneous materials is not limited to Cauchy continua. A robust homogenization framework based on higher-order continuum theories could enable the POD-TANN workflow to model complex microstructures that account for localization and bifurcation, thereby expanding the applicability of the approach.
5.2 Macroelement derivation, calibration and reuse
@@
An important consideration in this context is the number of POD modes used in the model, as this directly influences the dimensionality of the internal state variable vector. The selection of POD modes is crucial, as it determines how well the reduced-order model can capture the essential features of the system. For TL to be effective across different scenarios, the selected number of POD modes should be representative across a range of macro-model parameters, including variations in geometry, material properties, and boundary conditions, as stated before. This requirement could ensure the number of POD modes (i.e. of ISV vector components) to remain unchanged and sufficient for new scenarios, thereby enabling the successful application of TL, without TANN’s hyperparameters changes (e.g., the number of inputs, and consequently the number and density of hidden layers).
@@ 6 Conclusions
The new POD-TANN methodology integrates Proper Orthogonal Decomposition (POD) into the training of Thermodynamics-based Artificial Neural Networks (TANN, [7]) to model the macroscopic behavior of inelastic systems with microstructure.
The POD is applied to datasets of internal coordinates (IC) that describe microscale processes in heterogeneous, inelastic materials, to obtain the interal state variables (ISVs) at the macroscale. This method leverages the hierarchical structure of POD modes, ensuring that the first N modes capture the most significant information. The ISV vector’s components are hierarchically sorted, allowing dimensionality reduction to be decoupled from TANN training, offering flexibility in the number of ISV components used based on the desired accuracy.
One major advantage is the ability to incorporate microscopic information into the macroscale model using a simple and robust linear dimensionality reduction technique. The invertibility of the linear mapping enables both macroscopic homogenization and microscopic field reconstruction. Depending on the objective - whether macroscopic behavior or detailed microscopic fields description - the number of POD modes can be adjusted to balance training efficiency and reconstruction accuracy.
The methodology is demonstrated through three increasingly complex applications: homogenization of inelastic RUCs with continuous microstructure and a 3D geotechnical monopile model subjected to horizontal loading. Additionally, a RUC with spatially correlated random fields of constitutive parameters is investigated in the Appendix. In each case, the trained network accurately reproduced stress-strain or force-displacement paths, and the reconstructed microscopic fields’ errors were analyzed based on the number of POD modes.
The method excludes energy terms in the loss function during training, making it applicable to diverse data sources. These could include experimental data without energy measurements.
Overall, the POD-TANN approach offers a versatile and powerful tool for modeling heterogeneous inelastic microstructured systems, ensuring thermodynamic consistency and efficient model reduction. The results are promising and suggest the potential for even more complex applications.
@@ 7 Acknowledgment
The authors gratefully thank the funding of the European Research Council (ERC) under the Horizon 2020 research and innovation program of the European Union (Grant agreement ID 757848 CoQuake). In addition, the authors would like to thank Dr. Filippo Masi and the whole CoQuake group for our fruitful discussions.
@@
Appendix B: Performance on a Randomly Heterogeneous RUC with Spatially Correlated Constitutive Parameters
Examples presented in sections 4.1 and 4.2 feature single inclusions of increasing topological complexity. A RUC with a more complex, spatially correlated random material properties is presented in this appendix.
Correlated random field simulation algorithm
The generation of a correlated random field and its subsequent scaling to specified statistical properties involves a series of steps combining spatial domain considerations with transformations in the Fourier domain.
Initially, the domain is established as a three-dimensional grid characterized by dimensions $L_{x}$, $L_{y}$, and $L_{z}$ and corresponding resolutions $N_{x}$, $N_{y}$, and $N_{z}$. The spatial resolutions, given by $dx=\frac{L_{x}}{N_{x}}$, $dy=\frac{L_{y}}{N_{y}}$, and $dz=\frac{L_{z}}{N_{z}}$, determine the discrete spacing between points in the spatial field.
Wave numbers for each dimension are calculated using the formula $k_{x}=2\pi\cdot\text{FFT}(N_{x},d=dx)$, and similarly for $k_{y}$ and $k_{z}$, where FFT is the Fast Fourier Transform. These wave numbers are utilized to create a three-dimensional meshgrid, which serves as the foundation for the Fourier domain representation of the field. The radial wave number $k$ is computed as $k=\sqrt{\mathbf{k}_{x}^{2}+\mathbf{k}_{y}^{2}+\mathbf{k}_{z}^{2}}$, encapsulating the magnitude of the wave vector at each point in the Fourier space.
A Power Spectral Density (PSD) function $S(k)$ is defined to model the energy distribution across frequencies. In this application, a Gaussian decay model $S(k)=e^{-\frac{k^{2}}{\kappa^{2}}}$ is employed, where $\kappa$ modulates the spectral decay, influencing the smoothness and correlation properties of the spatial field. Figure B.15-a depicts the Gaussian form assumed in this study, with $\kappa=5$.
The procedure continues with the generation of complex Gaussian white noise $\eta(\mathbf{k})$ in the Fourier domain. The latter is then modulated by the square root of the PSD to form the Fourier transform of the random field, expressed as $\hat{f}(\mathbf{k})=\sqrt{S(\mathbf{k})}\cdot\eta(\mathbf{k})$. The modulated noise is transformed back to the spatial domain using the inverse Fast Fourier Transform (IFFT), resulting in the correlated random field $f(\mathbf{x})$. This field exhibits spatial correlations as prescribed by the PSD. To align the statistical properties of the field with specific requirements, the mean and standard deviation of $f(\mathbf{x})$ are first calculated. The field is then scaled from its original mean $\mu$ and standard deviation $\sigma$ to the desired mean $\mu^{\prime}$ and standard deviation $\sigma^{\prime}$ using the formula $f^{\prime}(\mathbf{x})=\frac{f(\mathbf{x})-\mu}{\sigma}\times\sigma^{\prime}+\mu^{\prime}$. This scaling ensures that the new field $f^{\prime}(\mathbf{x})$ adheres to the specified mean and variance, maintaining the correlation structure imparted by the PSD. This generation strategy has been exploited to ensure periodicity of the generated field, inherited by the Fourier operator, differently from other techniques, e.g., the ones based on Choleski decomposition, see [48].
Data generation
The RUC voxels have been modeled using a Drucker-Prager constitutive model, with an additional volumetric cap, as detailed in [33]. Spatially correlated random constitutive parameters have been assigned to the geometric field, which is derived from the coordinates of the Gauss points in the elements of the computational model.
Table 3 summarizes the statistical properties of the constitutive parameters used: $\beta$ and $d$ represent the friction angle and cohesion of the material, respectively. The parameter $K$ controls the dependence of the yield surface on the intermediate principal stress, while $E$ and $\nu$ are the elastic constants. The preconsolidation pressure is denoted by $p_{0}$, and the parameters $R$ and $\alpha$ further define the shape and hardening behavior of the cap. Figure B.1-b illustrates the spatial distribution of Young’s modulus assigned to the RUC realization. The dataset for training and validation was constructed by applying 20,000 random strain increments to the RUC, initially brought to a volumetric contraction of $0.1\%$. Further data augmentation was performed through dataset rotation, as described in Section 4.1.1.
$\beta$ (°) $d$ (kPa) $E$ (kPa) $\nu$ (-) $p_{0}$ (kPa) $\alpha$ (-) $R$ (-) $K$ (-)
Mean 40 15 18000 0.3 100 0.05 1.2 0.8
Std. 0.1 0.5 5000 0.01 0.5 - - -
Table 3: Mean and standard deviation of the used constitutive parameters for the elas tic-strain hardening plastic Drucker-Prager constitutive model with cap. Parameters without standard deviation are considered homogeneous.
Figure B.1: a) Gaussian decay used to spatially correlate the generated random field. b) Spatially correlated distribution of Young’s moduli in the RUC realization.
Figure B.2: a) SVD decomposition of the IC set. b) Training losses considering 10, 20 and 31 POD modes.
Training of the TANN
The training of the TANN energy network was guided by the analysis of the SVD of the internal coordinate dataset. Figure B.2-a presents both the normalized singular values and their cumulative sum. Three cases were considered for truncating the POD modes: 10 modes ($CR=99.9\%$), 20 modes ($CR=99.8\%$), and 31 modes ($CR=99.7\%$). The latter case corresponds to a significant drop in the relative magnitude of the singular values, marking the threshold for an almost complete representation of the original field. Figure B.2-b shows the learning curves for each case, reflecting the use of different numbers of POD modes. The hyperparameters for the energy network were inherited from those detailed in Section 4.1.2. Training was stopped after 2000 epochs, and in all cases, the loss decreased smoothly. The final MAE reached approximately 3e-3 for stresses and 5e-4 for the rate of dissipation, consistent with the loss definition provided in Section 3.3.
Inference of the TANN
Figure B.3: Comparison of the TANN inference capabilities along an unseen random path, considering 10, 20 and 31 POD modes retained.
The inference capabilities of the TANN are demonstrated on an unseen random strain path using the three trained models. Figure B.3 compares the simulated data with the TANN predictions. Although all three models reached a similar loss by the end of the training process, the accuracy in inference improves as the number of retained POD modes increases. Specifically, for the random strain path considered, the MAE is 3.21e-01 with 10 modes, 8.79e-02 with 20 modes, and 1.28e-02 with 31 modes.